# Impromptu Cybercrime Euphemism Detection

### v1

@@ Abstract
Xiang Li1,2, Yucheng Zhou3, Laiping Zhao4, Jing Li1, Fangming Liu1,2thanks:  Corresponding author. 1 Harbin Institute of Technology(Shenzhen), 2 Pengcheng Laboratory 3 SKL-IOTSC, CIS, University of Macau, 4 Tianjin University lixiang_alex@stu.hit.edu.cn, fangminghk@gmail.com
@@ 2.2 Context Modeling
Context modeling is a technique that enhances the understanding of contextual information in language models by training on specific objectives. Devlin et al. (2019) propose BERT that employs Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) to capture bidirectional contextual information, enhancing its expressive capabilities. These objectives enable BERT to capture bidirectional contextual information, thereby enhancing its expressive capabilities. XLNet, introduced by Yang et al. (2019), utilizes generalized auto-regressive learning to acquire bidirectional context, addressing pretraining and fine-tuning inconsistencies and overcoming the limitations of traditional auto-regressive models. To enhance BERT, Clark et al. (2020) propose ELECTRA that utilizes a replaced token detection (RTD) objective, allowing for a broader context and reducing redundancy in token prediction during pretraining. Moreover, Gao and Callan (2021) introduced Condenser, which incorporates an additional Transformer layer dedicated to extracting a condensed vector representation from the output of the language model. This condensed representation serves as the encoding for the textual sequence, enhancing the model’s ability to capture relevant contextual information.
@@ Acknowledgements
This work was supported in part by the Major Key Project of PCL under Grant PCL2024AS103 and PCL2022A05, and in part by the Shenzhen Science and Technology Program under Grant RCJC20231211085918010.

### v2

@@ Abstract
Xiang Li1,2, Yucheng Zhou3, Laiping Zhao4, Jing Li1222Corresponding Authors., Fangming Liu1,$2$Corresponding Authors. 1 Harbin Institute of Technology(Shenzhen), 2 Pengcheng Laboratory 3 SKL-IOTSC, CIS, University of Macau, 4 Tianjin University lixiang_alex@stu.hit.edu.cn, jingli.phd@hotmail.com, fangminghk@gmail.com
@@ 2.2 Context Modeling
Unlike feature augmentation Li and Zhou (2023), context modeling is a technique that enhances the understanding of contextual information in language models by training on specific objectives. Devlin et al. (2019) propose BERT that employs Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) to capture bidirectional contextual information, enhancing its expressive capabilities. These objectives enable BERT to capture bidirectional contextual information, thereby enhancing its expressive capabilities. XLNet, introduced by Yang et al. (2019), utilizes generalized auto-regressive learning to acquire bidirectional context, addressing pretraining and fine-tuning inconsistencies and overcoming the limitations of traditional auto-regressive models. To enhance BERT, Clark et al. (2020) propose ELECTRA that utilizes a replaced token detection (RTD) objective, allowing for a broader context and reducing redundancy in token prediction during pretraining. Moreover, Gao and Callan (2021) introduced Condenser, which incorporates an additional Transformer layer dedicated to extracting a condensed vector representation from the output of the language model. This condensed representation serves as the encoding for the textual sequence, enhancing the model’s ability to capture relevant contextual information.
@@ Acknowledgements
This work was supported in part by the Major Key Project of PCL under Grant PCL2024AS103 and PCL2022A05, the Shenzhen Science and Technology Program under Grant RCJC20231211085918010, National Science Foundation of China (62476070), Shenzhen College Stability Support Plan (GXWD20231128103232001) and Department of Science and Technology of Guangdong (2024A1515011540).